Prove that the sequence {an/n} is convergent where { an} is a bounded sequence


1
Expert's answer
2022-01-10T18:24:05-0500

As "a_n" is bounded, we know that there exists a positive real number "M" such that "|a_n|\leq M" for all "n\in \mathbb{N}". We will prove that the sequence "(\frac{a_n}{n})_{n\in\mathbb{N}}" converges to zero, let us fix "\varepsilon>0". We know that there exists "N\in\mathbb{N}" such that for all "n\geq N", "\frac{M}{n}< \varepsilon" (it is enough to take "N=\text{smallest integer greater than } M/\varepsilon"). We have then an estimate for any "n\geq N" :

"|\frac{a_n}{n}|\leq |\frac{M}{n}|<\varepsilon"

Therefore, "(a_n/n)_{n\in\mathbb{N}}\to 0".